# Bayesian Ensembling: Insights from Online Optimization and Empirical Bayes

### v1

@@ Appendix C Experimental Setup
C.1 Experimental Setup
All experiments were conducted using Ubuntu 22.04 with an Intel i9-9900K CPU with 128 GB of RAM and two NVIDIA Titan RTX GPUs. Code for our methods and experiments is available online with an MIT License.  [^2] https://github.com/DanWaxman/OnlineBayesianStacking, MIT License Implementations were in Jax/Objax (Bradbury et al., 2018; Objax Developers, 2020) [^3] https://github.com/google/objax, Apache 2.0 License based on modifying the codes of Waxman & Djurić (2024) [^4] https://github.com/DanWaxman/DynamicOnlineBasisExpansions, MIT License and Jones et al. (2024) [^5] https://github.com/petergchang/bong/tree/main, MIT License. Optimization algorithm implementations were adapted from and tested against the Universal Portfolios library (Vinkler, 2024) [^6] https://github.com/Marigold/universal-portfolios, MIT License.
C.2 Baselines
@@
The BCRP is the so-called “static” solution to our stacking problem, equivalent to the offline versions of Bayesian stacking presented in Hall & Mitchell (2007); Geweke & Amisano (2011). The resulting weights are a solution of Equation 6, i.e.,
@@
In regression problems without concept drift, this forms a reasonable baseline to measure regret with respect to.
@@
Our main baseline is O-BMA, which we consider to be the standard online ensembling approach in Bayesian applications. The weights are updated as Equation 3, i.e.,
@@
While DMA loses some of the nice statistical properties of O-BMA, it can be more robust to non-stationary scenarios, where the best expert make change over time. For the forgetting factor, we follow Raftery et al. (2010) and use a value of $\gamma=0.99$.
C.3 Portfolio Selection Algorithms
We include several OPS algorithms in our comparisons. We provide brief overviews of each, and our default set of hyperparameters, below.
@@ C.3 Portfolio Selection Algorithms
The EG algorithm (Helmbold et al., 1998) for portfolio selection follows the updates in Equation 10, i.e.,
@@
Weights are then obtained at each iteration by projection onto the simplex,
@@
Additionally, to help with very small market variability parameters, gradients may artificially be “smoothed” as
@@
The D-ONS algorithm (Yuan & Lamperski, 2020; Ding et al., 2021) is similar to the ONS algorithm (as in Hazan et al. (2007)), but uses a forgetting factor on the second-order information. We use the formulation of Ding et al. (2021), which the authors claim is more numerically stable than the discounting factor used in Yuan & Lamperski (2020). The D-ONS algorithm depends on parameters $\eta>0$ and $\gamma\in(0,1)$ and keeps track of a quantity
@@
Appendix D Details and Weights for the Subset of Linear Regressors Experiment
In this appendix, we provide more details on the experimental setup and results for the subset of linear regressions experiment in Section 5.1.
D.1 Experimental Details
@@
The OCO hyperparameters for OBS were set to reasonable default values and not tuned further:
@@ D.2 Weight Evolutions and Final Weights
As detailed in the main manuscript, one reason we may expect O-BMA to perform worse than OBS is the weight collapse of BMA. In this appendix, we empirically verify this property. In Figure 3, we show the final weights in the “open” subset linear regression experiment, for O-BMA, OBS (with EG and ONS algorithms), and the best constant rebalanced portfolio (BCRP) (i.e., the optimal pooling of Geweke & Amisano (2011), or Bayesian stacking from the prequential principle (Yao et al., 2018)). In Figure 4, we show the evolution of the BMA and OBS weights as new data arrives. Finally, in Figures 5 and 6, we show the analogous final weights and evolution of weights for the “closed” subset linear regression experiment.
Figure 3: The final weights in the “open” subset linear regression experiment.
Figure 4: The evolution of the weight vector $\boldsymbol{\mathbf{w}}_{t}$ as a function of $t$ in the “open” subset linear regression experiment. Results are shown for a single trial due to the noisy nature of the plots. Dots on the right side of a plot denote the final weights of the BCRP.
Figure 5: The final weights in the “closed” subset linear regression experiment.
Figure 6: The evolution of the weight vector $\boldsymbol{\mathbf{w}}_{t}$ as a function of $t$ in the “closed” subset linear regression experiment. Lines denote the median and shaded regions denote the 10th to 90th percentile over 10 trials. Dots on the right side of a plot denote the final weights of the BCRP.
Appendix E Details and Weights for the Online Variational Inference Experiment
In this appendix, we provide more details on the experimental setup and results for the online variational inference experiment in Section 5.2.
E.1 Experimental Details
As mentioned in the main text, we use the bong-dlr10-ef_lin model from Jones et al. (2024), which had the best performance out of variants tested in their experiments. The BONG performs variational inference using natural gradients, updating variational parameters $\boldsymbol{\mathbf{\psi}}_{t}$ as
where $\boldsymbol{\mathbf{F}}_{\boldsymbol{\mathbf{\psi}}_{t}}$ is the Fisher information matrix and $q_{\boldsymbol{\mathbf{\psi}}_{t}}$ is the variational posterior. The The bong-dlr10-ef_lin variant uses a variational family of Gaussians with low-rank covariance matrices (diagonal + a rank $10$ matrix), and approximates the predictive likelihood through a linearization using the empirical Fisher information matrix. We use the authors’ implementation of BONG. [^7] https://github.com/petergchang/bong/, MIT License
E.2 Weight Evolutions and Final Weights
Once again, we visualize the resulting weights in Figure 7 and Figure 8. We find a remarkable similarity between the final weights in all OCO algorithms tested and the BCRP solution, and that once again, BMA incorrectly collapses to a single model.
@@ E.2 Weight Evolutions and Final Weights
Appendix F Details and Weights for the Forecasting Experiment
In this appendix, we provide more details on the experimental setup and results for the forecasting experiment in Section 5.4.
F.1 Experimental Details
In the forecasting experiment, we use real data corresponding to the daily returns of the index S&P 500 from 2015 through 2020, consisting of 1257 unique observations. In econometrics, such data is often modeled with Generalized Autoregressive Conditional Hierarchical (GARCH) approaches, which provide probabilistic predictions for time series. We refer the read to Engle (2001) for a more detailed introduction to GARCH models.
@@ F.1 Experimental Details
In Geweke & Amisano (2011, 2012), the application of GARCH models to stock index data is also used to test the ensembling of probabilistic models. In GARCH, the conditional variance of the process is modeled as an autoregressive process depending on the lagged values of the conditional variance and the squared “innovations.” For instance, the GARCH(1,1)11(1,1)( 1 , 1 ) model with Gaussian innovations is given by:
$y_{t}$ $=\sigma_{t}\epsilon_{t},\enskip\epsilon_{t}\sim\mathcal{N}(0,\sigma_{t})$ (35)
@@
where $\epsilon_{t}$ are the innovations and $\sigma_{t}$ is the conditional variance of the process. For brevity, we only discuss the GARCH(1,1)11(1,1)( 1 , 1 ) with Gaussian innovations. The variants differ in the assumed process for the conditional variance.
We assign truncated Gaussian prior densities to $\boldsymbol{\theta}_{t}=[\alpha_{0},\alpha_{1},\beta,\sigma_{t}]$ and we are interested in sequential estimation of the posterior distribution $p(\boldsymbol{\theta}_{t}|y_{1:t})$. Sequential Monte Carlo (SMC) algorithms are very well suited for this task Doucet et al. (2001). An SMC algorithm recursively computes a particle approximation of the posterior,
@@
where $\boldsymbol{\theta}^{(i)}_{t}$ and $\rho_{t}^{(i)}$ denote respectively the particles and weights. After the arrival of $y_{t+1}$, the weights are recomputed and the particles propagated to form the particle approximation of the posterior at $t+1$. Using this particle approximation, we can obtain an approximation of the posterior predictive density,
@@
where $\boldsymbol{\theta}_{t+1}^{(i)}$ only differs from $\boldsymbol{\theta}^{(i)}_{t}$ in the component $\sigma^{(i)}_{t+1}$, that is obtained by substituting the previous set of particles in Eq. (36).
@@
where $w_{k,t}$ is the weight of model $k$ obtained at time $t$ with the approaches discussed in the paper. We run 10 independent simulations of this experiment considering 10 different random seeds.
F.2 Weight Evolutions and Final Weights
We visualize the resulting weights in Figure 9 and Figure 10.
@@ F.2 Weight Evolutions and Final Weights
Appendix G Details and Weights for the Non-Stationary Environment Experiment
@@
As mentioned in the main text, the models considered all belong to the dynamic online ensembles of basis expansions (DOEBE) family, which performs online GP regression using Kalman filtering and the random Fourier feature approximation Lázaro-Gredilla et al. (2010). Waxman & Djurić (2024) find that adding a random walk to the model parameters is important in capturing “dynamic” (i.e., nonstationary) behavior when applying approximate GPs to several real-world datasets, which can be related to “back-to-prior forgetting” (Van Vaerenbergh et al., 2012).
We create ensembles using an RFF GP with $100$ features, trained on the first 1000 data using the marginal likelihood. We then use a different value of the random walk scaled ($\sigma_{\text{rw}}\in\{10^{-4},10^{-3},10^{-2},10^{-1},10^{0}\}$), and ensemble them using O-BMA and various OBS algorithms.
In the interest of space, we only report the weight evolutions for SARCOS, Kuka #1, and Elevators in Figure 11. The other dataset results are qualitatively similar to Elevators.
(a) Elevators
(b) SARCOS
(c) Kuka #1
G.1 Cumulative Reward Plots
We show the cumulative reward plots (i.e., the average PLL as a function of $t$) in Figure 12.
@@ G.1 Cumulative Reward Plots
Appendix H Sensitivity to the Learning Rate
@@
For the experiment using the BONG (Section 5.2 & Appendix E), we performed additional experiments using differing hyperparameters in OCO algorithms. Specifically, we vary $\eta$ in EG and $\beta$ in ONS for values $10^{-k}$, for $k\in\{0,1,2,3,4\}$. Results can be found in Table 2. We find that the results are not particularly sensitive to the learning rate, with all results except for EG with high learning rate ($10^{0}$) and ONS with a small learning rate ($10^{-4}$) outperforming O-BMA.
@@
EG ($\eta$) $10^{0}$ $-1.012\pm 0.04$
$10^{-1}$ $-0.850\pm 0.02$
$10^{-2}$ $-0.850\pm 0.02$
$10^{-3}$ $-0.862\pm 0.02$
$10^{-4}$ $-0.865\pm 0.02$
ONS ($\beta)$ $10^{0}$ $-0.878\pm 0.03$
$10^{-1}$ $-0.859\pm 0.03$
$10^{-2}$ $-0.860\pm 0.02$
$10^{-3}$ $-0.900\pm 0.03$
$10^{-4}$ $-0.948\pm 0.03$
Soft-Bayes $-0.847\pm 0.02$
O-BMA $-0.932\pm 0.02$
BCRP $-0.843\pm 0.02$
Table 2: Comparison of the average predictive log likelihood value (higher is better) for the BONG experiment (Section 5.2) while varying the learning rates of EG and ONS. Results are reported as the median $\pm$ 1 standard deviation across five trials.
Results are similar in the forecasting example (Section 5.4 & Appendix F), where the results in Table 3 suggest largely comparable performance across learning rate values. The performance of EG degrades at very large or very small learning rates, but ONS seems fairly stable in this setting.
@@
Table 3: Comparison of the average predictive log likelihood value (higher is better) for the GARCH experiment (Section 5.4) while varying the learning rates of EG and ONS. Results are reported as the median $\pm$ 1 standard deviation across ten trials.

### v2

@@ Appendix B Experimental Setup
B.1 Experimental Setup
All experiments were conducted using Ubuntu 22.04 with an Intel i9-9900K CPU with 128 GB of RAM and two NVIDIA Titan RTX GPUs. Code for our methods and experiments is available online under an MIT License.  [^2] https://github.com/DanWaxman/OnlineBayesianStacking, MIT License Implementations were in Jax/Objax (Bradbury et al., 2018; Objax Developers, 2020) [^3] https://github.com/google/objax, Apache 2.0 License based on modifying the codes of Waxman & Djurić (2024) [^4] https://github.com/DanWaxman/DynamicOnlineBasisExpansions, MIT License and Jones et al. (2024) [^5] https://github.com/petergchang/bong/tree/main, MIT License. Optimization algorithm implementations were adapted from and tested against the Universal Portfolios library (Vinkler, 2024) [^6] https://github.com/Marigold/universal-portfolios, MIT License.
B.2 Baselines
@@
The BCRP is the so-called “static” solution to our stacking problem, equivalent to the offline versions of Bayesian stacking presented in Hall & Mitchell (2007); Geweke & Amisano (2011). The resulting weights are a solution to Equation˜6, i.e.,
@@
In regression problems without concept drift, this forms a reasonable baseline against which to measure regret.
@@
Our main baseline is O-BMA, which we consider to be the standard online ensembling approach in Bayesian applications. The weights are updated as Equation˜3, i.e.,
@@
While DMA loses some of the nice statistical properties of O-BMA, it can be more robust in non-stationary scenarios, where the best expert may change over time. For the forgetting factor, we follow Raftery et al. (2010) and use a value of $\gamma=0.99$.
B.3 Portfolio Selection Algorithms
We include several OPS algorithms in our comparisons. We provide brief overviews of each and our default set of hyperparameters, below.
@@ B.3 Portfolio Selection Algorithms
The EG algorithm (Helmbold et al., 1998) for portfolio selection follows the updates in Equation˜10, i.e.,
@@
Weights are then obtained at each iteration by projecting onto the simplex,
@@
Additionally, to help with very small market variability parameters, gradients may be artificially “smoothed” as
@@
The D-ONS algorithm (Yuan & Lamperski, 2020; Ding et al., 2021) is similar to the ONS algorithm (as in Hazan et al. (2007)), but it uses a forgetting factor for the second-order information. We use the formulation of Ding et al. (2021), which the authors claim is more numerically stable than the discounting factor used in Yuan & Lamperski (2020). The D-ONS algorithm depends on parameters $\eta>0$ and $\gamma\in(0,1)$ and keeps track of a quantity
@@
Appendix C Details and Weights for the Subset of Linear Regressors Experiment
In this appendix, we provide more details on the experimental setup and results for the subset of the linear regression experiment in Section˜5.1.
C.1 Experimental Details
@@
The OCO hyperparameters for OBS were set to reasonable default values and were not tuned further:
@@
Appendix D Details and Weights for the Online Variational Inference Experiment
In this appendix, we provide more details on the experimental setup and results for the online variational inference experiment in Section˜5.2.
D.1 Experimental Details
As mentioned in the main text, we use the bong-dlr10-ef_lin model from Jones et al. (2024), which had the best performance among the variants tested in their experiments. The BONG performs variational inference using natural gradients, updating variational parameters $\boldsymbol{\mathbf{\psi}}_{t}$ as
where $\boldsymbol{\mathbf{F}}_{\boldsymbol{\mathbf{\psi}}_{t}}$ is the Fisher information matrix and $q_{\boldsymbol{\mathbf{\psi}}_{t}}$ is the variational posterior. The bong-dlr10-ef_lin variant uses a variational family of Gaussians with low-rank covariance matrices (diagonal + a rank $10$ matrix) and approximates the predictive likelihood through a linearization using the empirical Fisher information matrix. We use the authors’ implementation of BONG. [^7] https://github.com/petergchang/bong/, MIT License
@@ D.2 Weight Evolutions and Final Weights
Once again, we visualize the resulting weights in Figure˜5 and Figure˜6. We find a remarkable similarity between the final weights in all OCO algorithms tested and the BCRP solution, and that once again, BMA incorrectly collapses to a single model.
@@
Appendix E Details and Weights for the Forecasting Experiment
In this appendix, we provide more details on the experimental setup and results for the forecasting experiment in Section˜5.3.
E.1 Experimental Details
In the forecasting experiment, we use real data corresponding to the daily returns of the S&P 500 index from 2015 through 2020, consisting of 1257 unique observations. In econometrics, such data are often modeled using Generalized Autoregressive Conditional Heteroskedasticity (GARCH) approaches, which provide probabilistic predictions for time series. We refer the reader to Engle (2001) for a more detailed introduction to GARCH models.
@@ E.1 Experimental Details
In Geweke & Amisano (2011; 2012), the application of GARCH models to stock index data is also used to test the ensembling of probabilistic models. In GARCH, the conditional variance of the process is modeled as an autoregressive process depending on the lagged values of the conditional variance and the squared “innovations.” For instance, the GARCH(1,1)(1,1) model with Gaussian innovations is given by:
$y_{t}$ $=\sigma_{t}\epsilon_{t},\enskip\epsilon_{t}\sim\mathcal{N}(0,1)$ (39)
@@
where $\epsilon_{t}$ are the innovations and $\sigma_{t}$ is the conditional variance of the process. For brevity, we only discuss the GARCH(1,1)(1,1) with Gaussian innovations. The variants differ in the assumed process for the conditional variance.
We assign truncated Gaussian prior densities to $\boldsymbol{\theta}_{t}=[\alpha_{0},\alpha_{1},\beta,\sigma_{t}]$, and we are interested in sequential estimation of the posterior distribution $p(\boldsymbol{\theta}_{t}|y_{1:t})$. Sequential Monte Carlo (SMC) algorithms are very well suited for this task Doucet et al. (2001). An SMC algorithm recursively computes a particle approximation of the posterior,
@@
where $\boldsymbol{\theta}^{(i)}_{t}$ and $\rho_{t}^{(i)}$ denote, respectively, the particles and weights. After the arrival of $y_{t+1}$, the weights are recomputed, and the particles are propagated to form the particle approximation of the posterior at $t+1$. Using this particle approximation, we can obtain an approximation of the posterior predictive density,
@@
where $\boldsymbol{\theta}_{t+1}^{(i)}$ only differs from $\boldsymbol{\theta}^{(i)}_{t}$ in the component $\sigma^{(i)}_{t+1}$, which is obtained by substituting the previous set of particles in Eq. (40).
@@
where $w_{k,t}$ is the weight of model $k$ obtained at time $t$ using the approaches discussed in the paper. We ran 10 independent simulations of this experiment, considering 10 different random seeds.
E.2 Weight Evolutions and Final Weights
We visualize the resulting weights in Figure˜7 and Figure˜8.
@@ E.2 Weight Evolutions and Final Weights
Appendix F Details and Weights for the Non-Stationary Environment Experiment
@@
As mentioned in the main text, the models considered all belong to the dynamic online ensembles of basis expansions (DOEBE) family, which perform online GP regression using Kalman filtering and the random Fourier feature approximation Lázaro-Gredilla et al. (2010). Waxman & Djurić (2024) find that adding a random walk to the model parameters is important in capturing “dynamic” (i.e., nonstationary) behavior when applying approximate GPs to several real-world datasets, which can be related to “back-to-prior forgetting” (Van Vaerenbergh et al., 2012).
We create ensembles using an RFF GP with $100$ features, trained on the first 1000 data points using the marginal likelihood. We then use different values of the random walk scale ($\sigma_{\text{rw}}\in\{10^{-4},10^{-3},10^{-2},10^{-1},10^{0}\}$) and ensemble them using O-BMA and various OBS algorithms.
In the interest of space, we only report the weight evolutions for SARCOS, Kuka #1, and Elevators in Figure˜10. The other dataset results are qualitatively similar to Elevators.
@@
(a) Elevators
(b) SARCOS
(c) Kuka #1
F.1 Cumulative Reward Plots
We show the cumulative reward plots (i.e., the average PLL as a function of $t$) in Figure˜9.
Appendix G Sensitivity to the Learning Rate
@@
For the experiment using the BONG (Section˜5.2 & Appendix˜D), we performed additional experiments using differing hyperparameters in OCO algorithms. Specifically, we vary $\eta$ in EG and $\beta$ in ONS for values $10^{-k}$ and $k\in\{0,1,2,3,4\}$. Results can be found in Table˜3. We find that the results are not particularly sensitive to the learning rate, with all results except for EG with a high learning rate ($10^{0}$) and ONS with a small learning rate ($10^{-4}$) outperforming O-BMA.
@@
EG ($\eta$) $10^{0}$ $1.012\pm 0.04$
$10^{-1}$ $0.850\pm 0.02$
$10^{-2}$ $0.850\pm 0.02$
$10^{-3}$ $0.862\pm 0.02$
$10^{-4}$ $0.865\pm 0.02$
ONS ($\beta)$ $10^{0}$ $0.878\pm 0.03$
$10^{-1}$ $0.859\pm 0.03$
$10^{-2}$ $0.860\pm 0.02$
$10^{-3}$ $0.900\pm 0.03$
$10^{-4}$ $0.948\pm 0.03$
Soft-Bayes $0.847\pm 0.02$
O-BMA $0.932\pm 0.02$
BCRP $0.843\pm 0.02$
Table 3: Comparison of the average predictive log likelihood value (higher is better) for the BONG experiment (Section˜5.2) while varying the learning rates of EG and ONS. Results are reported as the median $\pm$ 1 standard deviation across five trials.
Results are similar in the forecasting example (Section˜5.3 & Appendix˜E), where the results in Table˜4 suggest largely comparable performance across learning rate values. The performance of EG degrades at very large or very small learning rates, but ONS seems fairly stable in this setting.
@@
Table 4: Comparison of the average predictive log likelihood value (higher is better) for the GARCH experiment (Section˜5.3) while varying the learning rates of EG and ONS. Results are reported as the median $\pm$ 1 standard deviation across ten trials.
Appendix H Dynamic Model Averaging With Differing Forgetting Factors
In our experiments in Section˜5.4, we used a default “forgetting factor” for DMA, as recommended in Raftery et al. (2010). This is similar to our other experiments, in which we leave OBS hyperparameters at reasonable defaults and do not otherwise tune them. Nevertheless, to further understand the impacts of this hyperparameter, we repeated our experiments with more aggressive and more conservative values of the forgetting factor. We report results in Table˜5.
Overall, we find that DMA with extremely low forgetting factors performs well in highly nonstationary environments (Kuka #1, CaData), but it is still comparable to D-ONS, which has more robust guarantees regarding predictive performance. Moreover, this is without tuning the hyperparameters of D-ONS. In other datasets tested, OBS methods are still superior to DMA (CPU Small, SARCOS, Elevators). We additionally remark that DMA has much similarity with a sliding window approach and loses many of the statistical properties of O-BMA; on the other hand, adaptive/dynamic OCO algorithms D-ONS still come with robust theoretical guarantees.
Method Elevators SARCOS Kuka #1 CaData CPU Small
O-BMA -0.59 0.77 0.61 -0.84 0.33
DMA -0.59 0.79 0.69 -0.81 0.33
DMA-0.9 -0.60 0.80 0.75 -0.80 0.33
DMA-0.95 -0.59 0.80 0.73 -0.80 0.33
EG -0.58 0.78 0.56 -0.81 0.35
Soft-Bayes -0.58 0.80 0.70 -0.81 0.34
BCRP -0.57 0.80 0.66 -0.81 0.35
ONS -0.58 0.80 0.70 -0.80 0.34
D-ONS -0.58 0.81 0.73 -0.80 0.35
Table 5: Comparison of methods on non-stationary datasets (c.f. Table˜2), including DMA with additional forgetting factors.